OpenStudy (anonymous):

What is the vertex form of y=x^2-2x-3

OpenStudy (anonymous):

First, complete the square: \[y=x^{2}-2x+1-4\] and because \[(x^{2}-2x+1)=(x-1)^{2}\] \[y=(x-1)^{2}-4\] and that is the vertex form.

OpenStudy (anonymous):

Okay do you know how to take that same equation into intercept form???

OpenStudy (anonymous):

Factor \[x^{2}-2x-3=(x-3)(x+1)\] so, \[y=(x-3)(x+1)\] is the intercept form.

OpenStudy (anonymous):

Okay thank you so much i think i got it now.

OpenStudy (anonymous):

No problem! :)